OpenStudy (anonymous):

help with quadratic formula 2x^2 - 7x - 30 = 0

OpenStudy (anonymous):

\[x=-b \pm \sqrt{b ^{2}-4ac}/2a\]

hero (hero):

2x^2 - 7x-30 = 0 2x^2-12x+5x - 30 = 0 2x(x-6)+5(x-6) = 0 (x-6)(2x+5) = 0 x = 6,-5/2

OpenStudy (anonymous):

let 2= a , b= 7, c= -30 just put into the formula

OpenStudy (anonymous):

Hero is correct. Patissocoolike showed incorrect variables: let \[a=2 b=-7 c=-30\] the difference, of course, would be that b is -7 not 7

OpenStudy (anonymous):

so does the -7 become 7 (-b)

OpenStudy (anonymous):

correct, since b is -7 when the equation calls for a -b the correct input will be +7

OpenStudy (anonymous):

thx

hero (hero):

Just factor it

hero (hero):

No need for quad formulas if it's factorable

OpenStudy (anonymous):

it is -7 but in the formula you automatically know you must use the opposite sign of b